# Ionization and temperature measurements in warm dense copper using x-ray absorption spectroscopy

### v1

@@ Abstract
We detail experimental results inferring ionization and temperature for warm dense copper plasmas at several times solid density (15 to 25 g/cm3) and temperatures of 10 to 21 eV. Experiments performed at the OMEGA Laser Facility generate uniform warm dense matter conditions via symmetric shock compression of a buried copper layer. The plasma is probed with a laser-generated x-ray source to collect the K-shell x-ray absorption spectrum. Fitting bound-bound absorption contributions from constituent charge states of copper provides an estimated of approximately 4 to 7 for these warm dense copper plasmas. We find that these partially ionized plasmas have K-edge shifts of 12 to 30 eV and bound-bound resonance 1s3p absorption shifts of 4 to 26 eV with respect to the cold K-edge. This study provides necessary experimental data to improve ionization and opacity models in the warm dense matter regime.
@@ I Introduction
Accurate measurements of WDM are difficult due to several experimental challenges. First, generating WDM in the laboratory often includes significant temporal and spatial gradients that complicate analysis of the experimental observables. Incorporating gradients in the analysis of experimental data, while possible, increases the uncertainties in the inferred plasma conditions[7, 8], thus a design limiting gradients is key for high-quality measurements in this regime. Further, the high densities of WDM generally require x-rays to probe the plasma; these plasmas are also not hot enough to produce appreciable self-emission for x-ray diagnostic purposes. These challenges require diagnostics that use an external x-ray source such as x-ray radiography, x-ray Thomson scattering, and x-ray absorption spectroscopy (XAS), which have been employed in experiments to measure WDM properties [9, 10, 11, 12, 13]. In the case of these experimental methods, a large uniform volume is desirable for high-quality measurements, and in the case of x-ray radiography and x-ray Thomson scattering, a fit with a full physics model is required to infer plasma properties from the measured signals. A uniform WDM volume measured with XAS can provide a robust means to infer important properties of the plasma, such as temperature and ionization, without reliance on a full physics model to forward fit to the spectrum. Here, we present experimental results at the OMEGA Laser Facility[14] where a uniform warm dense Cu plasma was generated and measured using XAS at WDM conditions.
Our experiments use direct laser ablation to launch two counter-propagating shocks that heat and compress a buried layer of Cu to WDM conditions. The counter-propagating shocks generate a uniform volume of WDM for a short time after they have propagated completely through the buried layer and rebounded. The x-ray backlighter is timed to probe over this window and XAS is measured through the WDM sample. At these high-density conditions, the measured Cu K-edge exhibits sensitivity to the electron temperature, allowing for a direct inference of the temperature from the slope of the Cu K-edge. Additionally, the presence of bound-bound 1s3p resonance absorption is used to infer the charge state distribution of Cu. From simulation and spectral diagnostics, we infer the buried Cu foils reaches several times solid density (15 to 25 g/cm3) and temperatures ranging from 10 to 21 eV, and we find an average ionization of 4–7 across the achieved conditions. This study provides necessary experimental data to improve the predictive capabilities of models in the WDM regime.
@@ II Modeling Ionization of Warm Dense Copper
Ionization is a fundamental process in any plasma system and is especially important to understand in the WDM regime because the dense plasma environment modifies the embedded atoms’ bound states and consequently the ionization balance. Generally, the dense plasma lowers the effective ionization potential of the ions, due to increased screening from the surrounding high electron densities and lowering of the continuum due to close neighboring ions. This can impact the EOS through changes in ionization, and consequently other state variables such as temperature and pressure. Ionization also impacts the opacity, as the ionization states present in the plasma will have different absorption cross-sections with bound-bound and bound-free absorption at characteristic energies affecting radiation transport through the system. An accurate model of the EOS and opacity in this regime requires a detailed knowledge of the ionization processes as a function of density, pressure, and temperature. EOS and opacity calculations also depend on the atomic model used to account for the dense plasma environment. These models in the WDM regime cannot include all the relevant atomic states and truncations are made to make the calculations tractable. Experimental data is necessary to test the models and assumptions used in creating an EOS or opacity table in WDM. X-ray absorption spectra measured in these experiments exhibit strong K-shell absorption features that encode key information about the constituent ions in the plasma as well as the effects of the high densities on the resulting spectra. These results can be compared against model approaches that estimate the density effects in the plasma and to test different theoretical approaches.
@@
For a solid density sample at zero temperature, the absorption around the K-edge depends on the relevant charge state’s 1s binding energy, the chemical potential (sensitive to the electron degeneracy), and the ionization potential depression of the system. As the system is heated, excited states will open channels for bound-bound resonance absorption, notably the 1s3p transitions at several eV temperatures and 1s2p transitions at >100 eV temperatures for Cu. At warm temperatures (1–100 eV), the plasma’s charge state distribution is narrow, and each constituent ion will, when probed with external x-rays, contribute to the K-edge or resonance absorption at an identifiable energy. Increases in density can lead to additional ionization, and the ionization-dependent 1s binding energy will further shift the absorption features. Additionally, broadening effects of the observed absorption and emission lines from the surrounding plasma will affect the inferred width of each contribution to the charge state distribution and thus should be included, but this will not affect our ability to infer the average ionization .
@@
Calculating the 1snp transition energies typically begins with an isolated atom model, and the code used for this analysis was Flexible Atomic Code (FAC)[15], and a density-dependent model is added on top to account for the plasma environment. Simple models of the density-dependent plasma polarization shift (PPS)[16] predict that at electron densities of 1–2 cm-3, the 1s3p lines shift by a few to 10 eV, and 1s2p will shift a few eV. With consideration of these shifts, we make a direct inference of the constituent charge state distribution (CSD) of the Cu plasma.
Comparison with different models requires constraining the plasma conditions in phase space, both temperature and density are important considerations, paired with the CSD to fully assess the density effects in WDM. Measuring the K-edge with XAS can offer a powerful and robust means to infer temperature conditions within a WDM plasma independent of the material choice. Under these conditions, the ability for a probe x-ray to photoionize will be primarily shaped by the availability of free states around the K-edge energy. Fermi-Dirac statistics govern the degenerate electron population, and the probability of a vacancy in the continuum near the K-edge, , will be the dominant contributor to the energy profile of the K-edge[10]. Here is the electron occupancy, which follows a Fermi-Dirac distribution:
where E is the energy, is the chemical potential, T is the temperature, and is the Boltzmann constant. The Fermi energy is , where is the electron density, and is electron mass. Inference of temperature from the edge requires the material to be sufficiently degenerate; the Fermi energy is higher than the electron temperature. The K-edge slope reaches a maximal value of 4T at the inflection point of the edge, from which the temperature can be directly read independently of the energy position of the edge. For warm dense Cu at several times solid density, 15 to 25 g/cm3, the Fermi energy will range from 35 to 60 eV. As temperatures increase towards the decreasing degeneracy will lead to a less sensitive edge due to thermal broadening of the electron population and widening of the CSD. While temperature sensitivity of the K-edge can still be the dominant edge effect, in general as the temperature nears the Fermi energy the K-edge shape of the non-degenerate material becomes unsuitable as a temperature inference[11, 10].
@@
There have been efforts in recent decades to experimentally measure and characterize warm dense metals. Figure 1 shows some key experiments in this regime of WDM that investigate dense plasma effects. The previous experiments can be categorized based on their experimental method: XFEL experiments shown in Figure 1 from Ciricosta et al. [19, 20] isochorically heats targets to 100 to 200 eV at solid densities; Short pulse laser systems are also able to achieve extreme temperatures (500 eV) at solid density as demonstrated by Hoarty et al. [21]; Pulsed power experiments by Gomez and Hansen et al. [11] measured Fe in MagLIF liners[22] at Z Machine achieving 15 g/cm3 and 10 eV; long pulse laser systems can shock heat targets, experiments by Benuzzi-Mounaix et al. [23] show capabilities of 3x solid density and 8 eV temperatures. Conversely, ramp compression has been employed by Jiang et al. to compress Co to 1.5–2.1x solid density at lower temperatures < 1 eV; Finally, experiments by Döpper et al. [24] measured ionization of Be at extreme conditions ( 7–55 g/cm3, T110–160 eV) with capsule implosions on the NIF[25]. Importantly, amongst these experiments shown, no agreement exists on the IPD model that best describes the ionization in this regime of warm dense plasmas. The experiments presented in this study span an area of the phase space denser than shown with XFEL or short pulse laser experiments, and moderate temperatures that have little experimental data with a mid-Z metal. Our results provide additional data points that are, importantly, designed to be uniform plasmas and yield a valuable constraint to any model describing ionization in warm dense plasmas.
@@ III Target Design and Experimental Configuration
Experiments were conducted at the OMEGA Laser Facility, which is a 60 beam laser able to focus up to 30 kJ of 351 nm light over 1–3 on target. The experimental configuration consists of a planar target and a separate backlighter positioned 3 mm away. At the center of the planar buried layer target is a 10 m thick Cu disc (8.96 g/cm3 solid density) with a 500 m diameter. This is surrounded by an Au washer cut to match the Cu and acts as an x-ray block for the x-ray diagnostics. The Cu and Au lie between two symmetric plastic (CH) ablators, each with a thickness of 125 m. A diagram of the target package is shown in Figure LABEL:fig:omegaplatform(a). The backlighter is a 6 m thick Ge, coated onto a 100 m graphite foil. Additionally, the backlighter has shielding to prevent emission from the expanding laser plume from being in the line of sight of any diagnostic. The laser drive uses 14 beams on each side of the primary target, producing 3.4 to 5.4 kJ total laser energy on target per side, configured with a standard square pulse of either 1, 2, or 3 ns duration to achieve a variation of shock strengths that compress and heat the Cu sample. The achieved laser intensity is 161–770 TW/cm2 over the three pulse length configurations. The laser spot is smoothed by with distributed phase plates (DPPs) for each beam and smoothing by spectral dispersion (SSD) to reduce hot-spots in the drive and improve uniformity. The laser spot is chosen to overfill the target such that the buried layer sees a planar shock across its entire diameter, beginning from a uniform beam of 880 m diameter shown in Figure LABEL:fig:omegaplatform(b). The Ge backlighter is separately timed and irradiated with six additional beams, which deposits approximately 1.2 kJ total laser energy in a 500 ps square pulse at a spot diameter of 140 m.
@@
The 500 ps backlighter duration sets the timing window of our time-integrated diagnostics and is timed relative to the main laser drive to probe when the plasma achieves uniform conditions just after the symmetric shocks propagate through the buried sample. Figure LABEL:fig:hydraplots(a) depicts the three pulse length configurations used in this study. Dedicated experiments were completed with one-sided versions of the targets with an added quartz window to measure the shock breakout with VISAR[26] at the same laser intensities. Analysis of the experiments dedicated to measuring the shock timing with VISAR is further detailed below, in Section IV.1. We used the multi-physics radiation-hydrodynamics code HYDRA[27] to simulate both the one- and two-sided experiments. We determined constant laser power multipliers for the one-sided experiments that match the shock timing measured with VISAR, yielding the shock trajectories shown in LABEL:fig:hydraplots(b). These multipliers were then used to model the two-sided experiments to ensure proper backlighter timing and to predict the temperatures and densities achieved in the Cu layer. The HYDRA predictions for the average temperature and density over the Cu layer are shown in Figure LABEL:fig:hydraplots(c) and (d), respectively. The backlighter timing windows over which the XAS measurements are made for each pulse are shown as shaded regions in Figure LABEL:fig:hydraplots. The longer pulse durations provide more uniformity over the timing window, whereas the high-drive (1 ns pulse) is the least uniform over the measurement duration.
@@
Our primary diagnostic was XAS, obtained using the EFX spectrometer[28] used to measure the x-ray absorption spectrum over the backlit timing window. The EFX diagnostic was configured with a flat Si (111) Bragg crystal to measure the energy range of 6.3 to 11.4 keV. The XAS signal was recorded on image plate with a filter pack of Mn, Fe, and W, which provides fiducial K and L edges for fitting the spectral dispersion on each shot. The filter edges are used in combination with a ray-trace model using the XICSRT[29] code to constrain the energy dispersion relation to within 3.5 eV. A representative image of the EFX data collected on these experiments is shown in Figure LABEL:fig:omegaplatform(c).
@@ IV.1 Experimental VISAR measurements
Each experiment has two VISAR channels with different sensitivities; one leg was set with 33.66 m/ns/fringe, and the second with 13.538 m/ns/fringe, to resolve fringe jump ambiguities to determine the shock velocity time history as a function of position across the target measured within the quartz window. Additionally, a streaked optical pyrometer (SOP) is employed on the same line-of-sight, which gives another indication of shock breakout in the quartz. Figure 2(b)-(d) shows an example dataset from the 2ns pulse configuration. The shock breakout into the quartz window for both ablator with Au heat shield only and with Cu can be seen at 2.15 ns and 2.40 ns, respectively, by the shift in the fringes in 2(b) and 2(c). The line-imaging capability of the VISAR at OMEGA also allows for a determination of the shock planarity across the buried layer target. The inferred transverse shock profile for all three pulse configurations is shown in Figure 2(e); this is extracted from the discontinuity of the fringes transverse to the shock direction. This gives a quantitative estimate of how large the planar region of the shock is for the laser configuration. The Cu layer is then chosen to be well within this planar area to best maintain uniformity for the XAS measurements. Figure 2(e) shows that for these VISAR experiments that a target diameter of 350 µm or less will see little variation in the planarity of the shock. The two-sided experimental iterations used an increased spot size to allow for a 500 m diameter foil while maintaining the same planarity of the propagating shock and the Cu region.
The relationship between the laser intensity and the measured shock breakout is shown in Figure 2(f). The timing information is then used to fit the 1D HYDRA simulations with the scalar multiplier to the intensity. We found each drive configuration required its own multiplier to fit the VISAR timing, and these multipliers were used to predict the timing window for the two-sided symmetric drive experiments detailed in this study.

### v2

@@ Abstract
[…]
@@ I Introduction
Accurate measurements of WDM are difficult due to several experimental challenges. First, generating WDM in the laboratory often includes significant temporal and spatial gradients that complicate analysis of the experimental observables. Incorporating gradients in the analysis of experimental data, while possible, increases the uncertainties in the inferred plasma conditions[7, 8], thus a design limiting gradients is key for high-quality measurements in this regime. Further, the high densities of WDM generally require x-rays to spectroscopically probe the plasma; these plasmas are also not hot enough to produce appreciable self-emission for x-ray diagnostic purposes. These challenges require diagnostics that use an external x-ray source such as x-ray radiography, x-ray Thomson scattering, and x-ray absorption spectroscopy (XAS), which have been employed in experiments to measure WDM properties [9, 10, 11, 12]. In the case of these experimental methods, a large uniform volume is desirable for high-quality measurements, and in the case of x-ray radiography and x-ray Thomson scattering, a fit with a full physics model is required to infer plasma properties from the measured signals. A uniform WDM volume measured with XAS can provide a robust means to infer important properties of the plasma, such as temperature and ionization, without reliance on a full physics model to forward fit to the spectrum. Here, we present experimental results at the OMEGA Laser Facility[13] where a uniform warm dense Cu plasma was generated and measured using XAS at WDM conditions.
Our experiments use direct laser ablation to launch two counter-propagating shocks that heat and compress a buried layer of Cu to WDM conditions. The counter-propagating shocks generate a uniform volume of WDM for a short time after they have propagated completely through the buried layer and rebounded. Similar counter-propagating shock techniques have been demonstrated on other experimental platforms, although under different conditions and scientific objectives[14, 15, 16]. An x-ray backlighter is timed to probe the sample during this uniform phase, enabling x-ray absorption spectroscopy measurements of the WDM layer.
At these high-density conditions, existing atomic and plasma models have difficulty reproducing the observed absorption spectra, motivating the development of an approach to infer temperature and ionization by characterizing spectral features near the K-edge. Additionally, the presence of bound-bound 1s3p resonance absorption is used to infer the charge state distribution of Cu. From simulation and spectral diagnostics, we infer the buried Cu foils reaches several times solid density (15 to 25 g/cm3) and temperatures ranging from 10 to 21 eV, and we find an average ionization of 4–7 across the achieved conditions. This study provides necessary experimental data to improve the predictive capabilities of models in the WDM regime.
@@ II Modeling Ionization of Warm Dense Copper
Ionization is a fundamental process in any plasma system and is especially important to understand in the WDM regime because the dense plasma environment modifies the embedded atoms’ bound states and consequently the ionization balance. Generally, the dense plasma lowers the effective ionization potential of the ions, due to increased screening from the surrounding high electron densities and lowering of the continuum due to close neighboring ions. This can impact the EOS through changes in ionization, and consequently other state variables such as temperature and pressure. Ionization also impacts the opacity, as the ionization states present in the plasma will have different absorption cross-sections with bound-bound and bound-free absorption at characteristic energies affecting radiation transport through the system. An accurate model of the EOS and opacity in this regime requires a detailed knowledge of the ionization processes as a function of density, pressure, and temperature. EOS and opacity calculations depend on the atomic model used to account for the dense plasma environment. These models in the WDM regime cannot include all the relevant atomic states and truncations are made to make the calculations tractable. Experimental data is necessary to test the models and assumptions used in creating an EOS or opacity table in WDM. X-ray absorption spectra measured in these experiments exhibit strong K-shell absorption features that encode key information about the constituent ions in the plasma as well as the effects of the high densities on the resulting spectra. These results can be compared against model approaches that estimate the density effects in the plasma and to test different theoretical approaches.
@@
For a solid density sample at zero temperature, the absorption around the K-edge depends on the relevant charge state’s 1s binding energy, the chemical potential (sensitive to the electron degeneracy), and the ionization potential depression of the system. As the system is heated, excited states will open channels for bound-bound resonance absorption, notably the 1s3p transitions at several eV temperatures and 1s2p transitions at >100 eV temperatures for Cu. At warm temperatures (1–100 eV), the plasma’s charge state distribution is narrow, and each constituent ion will, when probed with external x-rays, contribute to the K-edge or resonance absorption at an identifiable energy. Increases in density can lead to additional ionization, and the ionization-dependent 1s binding energy will further shift the absorption features. Additionally, broadening effects of the observed absorption and emission lines from the surrounding plasma will affect on the inferred width of each contribution to the charge state distribution and thus should be included, but this will not affect our ability to infer the average ionization .
@@
Calculating the 1snp transition energies typically begins with an isolated atom model, and the code used for this analysis was Flexible Atomic Code (FAC)[17], and a density-dependent model is added on top to account for the plasma environment. Simple models of the density-dependent plasma polarization shift (PPS)[18] predict that at electron densities of , the 1s3p lines shift by a few to 10 eV, and 1s2p will shift a few eV. With consideration of these shifts, we make a direct inference of the constituent charge state distribution (CSD) of the Cu plasma.
Comparing experimental data with different models requires well-constrained plasma conditions in phase space, including temperature, density, and the (CSD), to accurately capture density-dependent effects in WDM. Measuring the K-edge with XAS can offer a powerful and robust means to infer temperature conditions within a WDM plasma independent of the material choice. The absorption spectrum at the K-edge is strongly characterized by the photoelectric coupling of a 1s electron to a vacant near or in the continuum. The photoabsorption cross section can be written as
where R is the electric dipole moment operator acting on the initial 1s state, , and selecting a p state among the final states, . Under degenerate conditions, Fermi-Dirac statistics govern the degenerate electron population, and the probability of a vacancy in the continuum near the K-edge, , will be the dominant contributor to the energy profile of the K-edge. Here is the electron occupancy, which follows a Fermi-Dirac distribution:
where E is the energy, is the chemical potential, T is the temperature, and is the Boltzmann constant. The Fermi energy is , where is the electron density, and is electron mass.
Inference of temperature from the edge requires the material to be sufficiently degenerate, such that the Fermi energy exceeds the electron temperature. Under this condition, the temperature can be directly read independently of the energy position of the edge. Potential contributions to the K-edge shape from the initial state are insignificant, as is largely unaffected in this WDM regime. Meanwhile, the final state implicitly contains a dependence the electronic density of states (DOS). Where here, the DOS is calculated from the square modulus of the final state integrated over space. So long as the profile of the DOS remains slowly varying over the K-edge region, the inference of temperature at the slope is reliable. Previous work by Dorchies et al. [19] has explored the sensitivity of the K-edge to the temperature of Al and has shown that for T > other effects such as a varying DOS or thermal broadening will affect the K-edge shape and the reliability for direct temperature readings. The K-edge will progressively underreport the temperature, signifying our inference as a lower bound of temperatures in conditions closer where and are of a comparable value. A further effect that could be expected due to the density of states is the presence of other absorption features, 1s3d, though we can expect these non-dipole transitions to be far less probable than the 1snp transitions.
@@
There have been efforts in recent decades to experimentally measure and characterize warm dense metals. Figure 1 shows some key experiments in this regime of WDM that investigate dense plasma effects. The previous experiments can be categorized based on their experimental method: XFEL experiments shown in Figure 1 from Ciricosta et al. [22, 23] isochorically heats targets to 100 to 200 eV at solid densities; Short pulse laser systems are also able to achieve extreme temperatures (500 eV) at solid density as demonstrated by Hoarty et al. [24]; Pulsed power experiments by Gomez and Hansen et al. [10] measured Fe in MagLIF liners[25] at Z Machine achieving 15 g/cm3 and 10 eV; long pulse laser systems can shock heat targets, experiments by Benuzzi-Mounaix et al. [26] show capabilities of 3x solid density and 8 eV temperatures. Conversely, ramp compression has been employed by Jiang et al. to compress Co to 1.5–2.1x solid density at lower temperatures < 1 eV; Finally, experiments by Döpper et al. [27] measured ionization of Be at extreme conditions ( 7–55 g/cm3, T110–160 eV) with capsule implosions on the NIF[28]. Importantly, among these experiments shown, no agreement exists on the IPD model that best describes the ionization in this regime of warm dense plasmas. The experiments presented in this study span an area of the phase space denser than shown with XFEL or short pulse laser experiments, and moderate temperatures that have little experimental data with a mid-Z metal. Our results provide additional data points that are, importantly, designed to be uniform plasmas and yield a valuable constraint to any model describing ionization in warm dense plasmas.
@@ III Target Design and Experimental Configuration
Experiments were conducted at the OMEGA Laser Facility, which is a 60 beam laser able to focus up to 30 kJ of 351 nm light over 1–3 ns on target. The experimental configuration consists of a planar target and a separate backlighter positioned 3 mm away. At the center of the planar buried layer target is a 10 m thick Cu disc (8.96 g/cm3 solid density) with a 500 m diameter. This is surrounded by an Au washer cut to match the Cu and acts as an x-ray block for the x-ray diagnostics. The Cu and Au lie between two symmetric plastic (CH) ablators, each with a thickness of 125 m. A diagram of the target package is shown in Figure 2(a). The backlighter is a 6 m thick Ge, coated onto a 100 m graphite foil. Additionally, the backlighter has shielding to prevent emission from the expanding laser plume from being in the line of sight of any diagnostic. The laser drive uses 14 beams on each side of the primary target, producing 3.4 to 5.4 kJ total laser energy on target per side, configured with a standard square pulse of either 1, 2, or 3 ns duration to achieve a variation of shock strengths that compress and heat the Cu sample. The achieved laser intensity is 161–770 TW/cm2 over the three pulse length configurations. The laser spot is smoothed by with distributed phase plates (DPPs) for each beam and smoothing by spectral dispersion (SSD) to reduce hot-spots in the drive and improve uniformity. The laser spot is chosen to overfill the target such that the buried layer sees a planar shock across its entire diameter, beginning from a uniform beam of 880 m diameter shown in Figure 2(b). The Ge backlighter is separately timed and irradiated with six additional beams, which deposits approximately 1.2 kJ total laser energy in a 500 ps square pulse at a spot diameter of 140 m.
@@
The 500 ps backlighter duration sets the timing window of our time-integrated diagnostics and is timed relative to the main laser drive to probe when the plasma achieves uniform conditions just after the symmetric shocks propagate through the buried sample. Figure 3(a) depicts the three pulse length configurations used in this study. Dedicated experiments were completed with one-sided versions of the targets with an added quartz window to measure the shock breakout with VISAR[29] at the same laser intensities. Analysis of the experiments dedicated to measuring the shock timing with VISAR is further detailed below, in Section IV.1. We used the multi-physics radiation-hydrodynamics code HYDRA[30] to simulate both the one- and two-sided experiments. We determined constant laser power multipliers for the one-sided experiments that match the shock timing measured with VISAR, yielding the shock trajectories shown in Figure 3(b). These multipliers were then used to achieve the same laser intensities (per side) on the two-sided HYDRA models to ensure proper backlighter timing and to predict the temperatures and densities achieved in the Cu layer. The HYDRA predictions for the average temperature and density over the Cu layer are shown in Figure 3(c) and (d), respectively. The backlighter timing windows over which the XAS measurements are made for each pulse are shown as shaded regions in Figure 3. An accurately timed backlighter ensures a more uniform plasma for the measured XAS and for the inference of temperature from the Cu K-edge. The longer pulse durations provide more uniformity over the timing window, whereas the high-drive (1 ns pulse) is the least uniform over the measurement duration.
@@
Our primary diagnostic was XAS, obtained using the EFX spectrometer[31], which measures the K-edge x-ray absorption spectrum over the backlighter timing window. The EFX diagnostic was configured with a flat Si (111) Bragg crystal to cover an energy range from 6.3 to 11.4 keV. The XAS signal was recorded on image plate using a filter pack consisting of Mn, Fe, and W, which provides fiducial K- and L-edge features for fitting the spectral dispersion on each shot. These filter edges are used in combination with a ray-trace model implemented in the XICSRT code[32] to constrain the energy dispersion relation to within 3.5 eV.
For the backlighter source size used in these experiments (140 m diameter), we achieve a spectral resolution of approximately 10 eV at 9 keV (900), with source-size broadening providing the dominant limitation. Appendix A details the theoretical framework used to estimate the spectral resolution for this experimental configuration. In practice, changes in the K-edge slope corresponding to temperature differences greater than 3 eV are resolvable, and the minimum temperature distinguishable from the cold case is approximately 2 eV. A representative EFX image from these experiments is shown in Figure 2(c).
@@ IV.1 Experimental VISAR measurements
Each experiment has two VISAR channels with different sensitivities; one leg was set with 33.66 m/ns/fringe, and the second with 13.538 m/ns/fringe, to resolve fringe jump ambiguities to determine the shock velocity time history as a function of position across the target measured within the quartz window. Additionally, a streaked optical pyrometer (SOP) is employed on the same line-of-sight, which gives another indication of shock breakout in the quartz. Figure 4(b) shows an example dataset from the 2ns pulse configuration. The shock breakout into the quartz window for both ablator with Au heat shield only and with Cu can be seen at 2.15 ns and 2.40 ns, respectively, by the shift in the fringes in 4(b) and 4(c). The line-imaging capability of the VISAR at OMEGA also allows for a determination of the shock planarity across the buried layer target. The inferred transverse shock profile for all three pulse configurations is shown in Figure 4(c); this is extracted from the discontinuity of the fringes transverse to the shock direction. This gives a quantitative estimate of how large the planar region of the shock is for the laser configuration. The Cu layer is then chosen to be well within this planar area to best maintain uniformity for the XAS measurements. Figure 4(e) shows that for these VISAR experiments that a target diameter of 350 µm or less will see little variation in the planarity of the shock. The two-sided experimental iterations used an increased spot size to allow for a 500 m diameter foil while maintaining the same planarity of the propagating shock and the Cu region.
The relationship between the laser intensity and the measured shock breakout is shown in Figure 4(f). The timing information is then used to fit the 1D HYDRA simulations with the scalar multiplier to the intensity. We found each drive configuration required its own multiplier to fit the VISAR timing, and these multipliers were used to predict the timing window for the two-sided symmetric drive experiments detailed in this study. The multipliers needed to scale HYDRA to match the measured shock timing were 0.75, 0.56, and 0.41 for the low-, mid-, and high-drive configurations, respectively, and these calibrated simulations directly informed the timing of the backlighter in the primary XAS experiments.